OpenStudy (anonymous):

Differentiate the trigonometric function 1-sinx/cosx

OpenStudy (anonymous):

(1-sin(x))/cos(x) ?

OpenStudy (anonymous):

yep thats right

OpenStudy (anonymous):

you can avoid using quotient rule

OpenStudy (anonymous):

i used the quotient rule but my answer didnt come out right

OpenStudy (anonymous):

1/cos(x) - sin(x)/cos(x) D'(sec(x) - tan(x)) sec(x)tan(x) - sec(x)^2

OpenStudy (anonymous):

\[\frac{1}{cos(x)+1}\] yes?

OpenStudy (anonymous):

yeah thats right can u explain

OpenStudy (anonymous):

so many forms this can take

OpenStudy (anonymous):

let me write it out. may take a minute

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

wow i had it and i lost it. let me try this

OpenStudy (anonymous):

just use the quotient rule.. the denominator is \[cos^2(x)\] the numerator is \[-cos^2(x) + sin(x)(1-sin(x))\]

OpenStudy (anonymous):

the numerator simplifies to \[-cos^2(x) + sin(x)-sin^2(x)=-cos^2(x)-sin^2(x)+sin(x)\]

OpenStudy (anonymous):

=\[sin(x)-1\]

OpenStudy (anonymous):

so we get \[\frac{sin(x)-1}{cos^2(x)}\]

OpenStudy (anonymous):

oh i see i got the same except my sin^2(x) was positive aswell

OpenStudy (anonymous):

now i think we can rewrite the denominator

OpenStudy (anonymous):

no it should be negative i think.

OpenStudy (anonymous):

ok im gona have to check my working

OpenStudy (anonymous):

put \[f(x)=1-sin(x)\] so \[f'(x)=-cos(x)\] \[g(x)=cos(x)\] \[g'(x)=-sin(x)\] numerator is \[gf'-fg'=cos(x)(-cos(x))-(1-sin(x))(-sin(x))\]

OpenStudy (anonymous):

=\[-cos^2(x)+sin(x)(1-sin(x))\] \[=-cos^2(x)+sin(x)-sin^2(x)=sin(x)-1\]

OpenStudy (anonymous):

so at least we got that far. deminator is \[cos^2(x)=1-sin^2(x)=(1-sin(x))(1+sin(x))\]

OpenStudy (anonymous):

damn now i am getting \[\frac{-1}{1+sin(x)}\]

OpenStudy (anonymous):

maybe i screwed up somewhere

OpenStudy (anonymous):

no thats right

OpenStudy (anonymous):

i looked at the wrong answer before lol

OpenStudy (anonymous):

oh lord. what is the right answer in the book?

OpenStudy (anonymous):

yes it was lol

OpenStudy (anonymous):

what is it?

OpenStudy (anonymous):

-1/1+sin(x)

OpenStudy (anonymous):

whew!

OpenStudy (anonymous):

done finito halas stick a fork in it

OpenStudy (anonymous):

you still got it

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

:)